# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v1

@@ Abstract.
High-Performance Computing (HPC) systems excel in managing distributed workloads, and the growing interest in Artificial Intelligence (AI) has resulted in a surge in demand for faster methods of Machine Learning (ML) model training and inference. In the past, research on HPC I/O focused on optimizing the underlying storage system for modeling and simulation applications and checkpointing the results, causing writes to be the dominant I/O operation. These applications typically access large portions of the data written by simulations or experiments. ML workloads, in contrast, perform small I/O reads spread across a large number of random files. This shift of I/O access patterns poses several challenges to HPC storage systems. In this paper, we survey I/O in ML applications on HPC systems, and target literature within a -year time window from to . We provide an overview of the common phases of ML, review available profilers and benchmarks, examine the I/O patterns encountered during ML training, explore I/O optimizations utilized in modern ML frameworks and proposed in recent literature, and lastly, present gaps requiring further R&D. We seek to summarize the common practices used in accessing data by ML applications and expose research gaps that could spawn further R&D.
@@ 1. Introduction
Existing High-Performance Computing (HPC) Input and Output (I/O) research has focused on techniques to optimize performance when running traditional HPC application workloads, which typically include simulations and checkpointing partial results (Párraga et al., 2021). Because of the increased popularity of Machine Learning (ML) workloads, there is a rising demand for I/O systems that can effectively accommodate their distinct I/O access patterns. Write operation bursts commonly dominate traditional workloads; however, ML workloads are usually read-intensive and use many small files (Paul et al., 2020). Due to the absence of a well-established consensus on the preferred I/O stack for ML workloads, numerous developers resort to crafting their own ad-hoc algorithms and storage systems to cater to the specific requirements of their applications (Gainaru et al., 2022). This can result in sub-optimal application performance due to the under-utilization of the storage system, prompting the necessity for novel I/O optimization methods tailored to the demands of ML workloads.
@@
In Fig. 1, we show the evolving I/O stack used for running ML workloads (on the right side) in comparison with the traditional HPC I/O stack (on the left side). Traditional HPC I/O stack has been developed to support massive parallelism. Applications use the MPI programming model and I/O middleware, such as HDF5 and netCDF, which internally use MPI-IO and POSIX-IO interfaces to move (read or write) data between main memory and parallel file systems. While many ML workloads still use the traditional I/O stack on HPC systems, some that use remote cloud computing systems utilize data warehouses to store and retrieve large datasets. When using the traditional I/O stack, high bandwidth and low latency data access are possible due to data locality and the highly optimized stack. In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding and Taylor, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024). Given the high latency inherent in accessing data stored in a remote storage location, optimizing data retrieval is crucial for reasonable performance. Despite the lower I/O performance imposed when using a data warehouse, they are still preferred in many scenarios because they provide a centralized data storage location, facilitate data integration from various sources, and enable data governance, which can result in higher quality data. High-quality data is essential when training ML models because it directly influences their accuracy, reliability, and generalizability (Jain et al., 2020). In addition, the monetary cost needed to purchase and maintain an HPC system can far outweigh the cost of using managed cloud services (Carlyle et al., 2010; Marathe et al., 2013).
@@
The differences listed above in the I/O stacks for ML applications using the traditional PFS compared to cloud computing systems motivates the development of novel storage system designs and I/O performance optimization techniques to adapt to the I/O operations commonly found in ML processes. Ideally, the methods would be transparent to the application developers. Notably, many scientists may lack formal training in software development, as evidenced by Paul et al. (Paul et al., 2021). The study revealed that ML workloads are becoming ever more popular and that developers from various scientific domains beyond computer science often overlook the utilization of node-local storage (i.e., burst buffers) within HPC systems. Further evidence for the increasing interest in distributed training is the number of posts on the popular tech website StackOverflow. The number of posts discussing distributed training staggeringly increased from posts in to posts in 2021 (Liu et al., 2023a). The increasing number of ML workloads being run on HPC systems prompts a need for transparent I/O optimization techniques to decrease the complexity required for scientists to write software that takes advantage of the resources offered by HPC systems.
@@
Furthermore, scientists, I/O performance enthusiasts, and HPC practitioners can gain significant insights from benchmarks that emulate ML workloads accurately and profilers that capture critical I/O information from the underlying storage systems. By using these tools, individuals can find potential I/O bottlenecks within their ML applications. Access to realistic benchmarking tools and fine-grained profilers is also invaluable for HPC storage research, allowing researchers to analyze and emulate the diverse range of I/O operations found in ML workloads. Currently, I/O analysis tools often fall short in a number of ways, including the inability to identify I/O requests at the thread level and only capturing traces of specific I/O interfaces (i.e., POSIX and STDIO). This makes it challenging to find a general-purpose tool applicable to a wide variety of workloads, which can reduce the consistency in which I/O traces are captured. In addition, many tools (i.e., Darshan (Snyder et al., 2016a) and Scalasca (Geimer et al., 2010)) target MPI applications limiting their compatibility and effectiveness with other communication libraries (i.e., GLOO (Facebook Incubator, 2024), Spark (Apache Software Foundation, 2024), and Dask (Dask, 2024)). It should be noted that Darshan is capable of capturing I/O traces from applications that do not use MPI through the NON_MPI environment variable. However, this results in a single DXT log being output for each process, instead of a single file capturing all I/O traces when used to capture traces from an application using MPI. In some cases, this leads developers with the task of manually merging the results from their non-MPI applications in order to conduct further analysis.
@@
In this paper, our contributions are a presentation and analysis of the I/O patterns found when emulating two renowned ML workloads: BERT (Devlin et al., 2019a), and Unet3D (Özgün Çiçek et al., 2016). In addition, we conduct a review of the currently proposed I/O optimization techniques through R&D efforts and present several areas where future research is needed. Understanding the I/O patterns found in ML applications allows model training speeds to be reduced. This contributes to the rapid development of ML models, which can be used to perform tasks such as medical imaging analysis and help in the creation of autonomous vehicles. Analyzing existing R&D and identifying gaps in I/O research allows ML development to continue at its fast-growing pace while avoiding I/O bottlenecks. To ensure a comprehensive review of the current literature, we have selected a -year time window from to and used IEEE Xplore and ACM Digital Library to filter by papers that mention the following keywords: “I/O Access Patterns,” “I/O pattern,” “HPC,” “Storage performance,” and “I/O performance” in AI/ML applications. Additional papers were found through the citations of the work initially found through the digital libraries for a total of papers. The various subjects explored in the papers surveyed are shown in Table 1. We selected papers from leading conferences and journals including:
@@
IEEE International Symposium on Modeling, Analysis, and Simulation of Computer and Telecommunication Systems (MASCOTS)
@@
The structure of the paper is as follows. First, in Section 2, we describe the data formats and modalities found in ML workloads and provide an overview of the typical phases of ML applications, and the popular methods of distributing the model training on HPC systems. Next, in Section 3, we describe the benchmarks and profiling tools scientists and researchers can use to ensure proper system utilization under ML workloads. In Section 4, we describe the spatial and temporal I/O access patterns found in the training phase of ML applications by analyzing currently available representative I/O benchmarks. We then discuss, in Section 5, various I/O optimization techniques used in modern ML frameworks and those proposed in recent research literature. Lastly, in Section 6, we discuss the gaps in the current research of ML I/O on HPC systems.
2. ML Data: Data Formats, Modalities and Common Phases
In this section, we describe the ML lifecycle and the data formats and modalities used throughout. We also discuss the common phases ML applications undergo and ways of distributing these phases to exploit the resources offered by HPC systems. Finally, model checkpointing and its impact on training speeds are described.
ML applications typically undergo four phases during the ML lifecycle: data generation, dataset preparation, training, and inference (Párraga et al., 2021). We show a taxonomy of the data management operations and characteristics throughout this lifecycle in Fig. 2. During data generation, data is obtained through various sources and methods such as simulations, web scraping, and public datasets. During this phase, properties such as the dataset modality, file format, and data storage location are often decided. The data preparation phase is split into various data preparation steps, including transformation, augmentation, and splitting. The training arm of the taxonomy consists of data I/O operations across common training algorithms and checkpointing tasks. Inference includes a similar taxonomy as the training phase. As can be seen, data is a prominent component throughout the entirety of the ML lifecycle.
@@ 2.1. Data Formats and Dataset Modalities
Data is at the forefront of the ML lifecycle, and without the correct quantity and quality of data, machine learning models fail to achieve optimal performance and generalization, disabling their ability to provide accurate insights and solutions. Here, we describe the file formats and data modalities commonly used during model training and inference.
@@
LibriSpeech (Panayotov et al., 2015a) | GiB | .flac, .txt | ||
@@
UCR Time Series Archive (Dau et al., 2018) | MiB | .tsv | ||
@@
Text | refinedweb (Penedo et al., 2023) | TiB | .parquet | |
@@
The Stack v2 (Lozhkov et al., 2024) | TiB | Code Files |
The selection of file formats model developers can choose from is often constrained by the data modality in use. The dataset modality refers to the mode or type of data being stored (e.g., audio, video, image, etc.). Due to the large variety of ML model goals (e.g., image recognition, optical character recognition (OCR), relationship detection, etc.), modalities vary widely. The dataset modality has a significant correlation to the file format that is used and, therefore, the common I/O operations required to load and use the dataset. In Table 3, we list the dataset modalities commonly used for ML applications and three corresponding publicly available datasets often used for ML research, exemplifying each modality. The provided values are approximate and subject to variation depending on which version of the dataset is selected. For instance, the Charades (Sigurdsson et al., 2016) dataset is available at different video resolutions, while the CIFAR-10 (Krizhevsky, 2009) dataset is available as a custom binary format or as a binary containing a Python “Pickled” object. Additionally, in certain cases, like AudioSet, links to the video data are provided rather than the raw video data itself, allowing users to choose the preferred video format.

### v2

@@ Abstract.
Growing interest in Artificial Intelligence (AI) has resulted in a surge in demand for faster methods of Machine Learning (ML) model training and inference. This demand for speed has prompted the use of high performance computing (HPC) systems that excel in managing distributed workloads. Because data is the main fuel for AI applications, the performance of the storage and I/O subsystem of HPC systems is critical. In the past, HPC applications accessed large portions of data written by simulations or experiments or ingested data for visualizations or analysis tasks. ML workloads perform small reads spread across a large number of random files. This shift of I/O access patterns poses several challenges to modern parallel storage systems. In this paper, we survey I/O in ML applications on HPC systems, and target literature within a -year time window from to . We define the scope of the survey, provide an overview of the common phases of ML, review available profilers and benchmarks, examine the I/O patterns encountered during offline data preparation, training, and inference, and explore I/O optimizations utilized in modern ML frameworks and proposed in recent literature. Lastly, we seek to expose research gaps that could spawn further R&D.
@@ 1. Introduction
Existing High-Performance Computing (HPC) Input and Output (I/O) research has focused on techniques to optimize performance when running traditional HPC application workloads, which typically include simulations and checkpointing partial results (et al, 2021e). Because of the increased popularity of Machine Learning (ML) workloads, there is a rising demand for I/O systems that can effectively accommodate their distinct I/O access patterns. Write operation bursts commonly dominate traditional workloads; however, ML workloads are usually read-intensive and use many small files (et al, 2020f). Due to the absence of a well-established consensus on the preferred I/O stack for ML workloads, numerous developers resort to crafting their own ad hoc algorithms and storage systems to accommodate the specific requirements of their applications (Gainaru et al., 2022). This can result in suboptimal application performance due to the under-utilization of the storage system, prompting the necessity for novel I/O optimization methods tailored to the demands of ML workloads.
In Fig. 1, we show the I/O stack used to run ML workloads (on the right side) compared to the traditional HPC I/O stack (on the left side) (et al, 2021e). The traditional HPC I/O stack has been developed to support massive parallelism. Applications are typically written in C/C++/Fortran languages that use the MPI programming model and I/O middleware, such as HDF5 (The HDF Group, 2022; Byna et al., 2020) and netCDF (Lee et al., 2008), which internally use MPI-IO and POSIX-IO interfaces to move (read or write) data between main memory and parallel file systems (PFS). In contrast, ML workloads are typically written in productivity-oriented languages, such as Python, and use PyTorch and TensorFlow frameworks.
@@
Although many ML workloads still use the traditional I/O stack on HPC systems, some are shifting to use remote cloud computing systems that offer services such as data warehouses to store and retrieve large datasets. Cloud platforms such as Google Cloud Platform, AWS, and Microsoft Azure offer access to large datasets, including Google Analytics data (Google Cloud, 2024a) hosted on BigQuery (Google Cloud, 2024b), the Genomes Project (Amazon Web Services, 2024a) hosted on Amazon Redshift (Amazon Web Services, 2024b), and NYC Taxi & Limousine Commission data (Microsoft Azure, 2024b) hosted on Azure Synapse Analytics (Microsoft Azure, 2024a). HPC workloads that aim to use these datasets require near-first-class integration with the cloud to enable efficient data transfers to HPC systems. When using the traditional I/O stack, high bandwidth and low latency data access are possible due to the locality of the data and the highly optimized stack. In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding and Taylor, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024). Given the high latency inherent in accessing data stored in a remote storage location, optimizing data retrieval is crucial for reasonable performance. Despite the lower I/O performance imposed when using a data warehouse, they are still preferred in many scenarios because they provide a centralized data storage location, facilitate data integration from various sources, and enable data governance, which can result in higher-quality data. High-quality data is essential when training ML models because it directly influences their accuracy, reliability, and generalizability (et al, 2020c). In addition, the monetary cost needed to purchase and maintain an HPC system can far outweigh the cost of using managed cloud services (Carlyle et al., 2010; et al, 2013). Furthermore, integration of HPC systems with available cloud resources is necessary due to the cloud being used as a data storage solution from which datasets can be publicly shared and used by all.
@@
The differences listed above in the I/O stacks for ML applications using the traditional PFS compared to cloud computing systems motivate the development of novel storage system designs and I/O performance optimization techniques to adapt to the I/O operations commonly found in ML processes. Ideally, the methods would be transparent to the application developers. Notably, many scientists may lack formal training in software development, as evidenced by Paul et al. (et al, 2021g). The study revealed that ML workloads are becoming ever more popular and that developers from various scientific domains beyond computer science often overlook the utilization of node-local storage (i.e., burst buffers) within HPC systems. Further evidence for the increasing interest in distributed training is the number of posts on the popular tech website StackOverflow. The cumulative number of posts discussing distributed training increased staggeringly from posts in to posts in 2024 (Community, 2024). The increasing number of ML workloads that are running on HPC systems prompts the need for transparent I/O optimization techniques to decrease the complexity required for scientists to write software that takes advantage of the resources offered by HPC systems.
@@
Furthermore, scientists, I/O performance enthusiasts, and HPC practitioners can gain significant insights from benchmarks that accurately emulate ML workloads and profilers that capture critical I/O information from the underlying storage systems. Using these tools, individuals can find potential I/O bottlenecks within their ML applications. Access to realistic benchmarking tools and fine-grained profilers is also invaluable for HPC storage research, allowing researchers to analyze and emulate the diverse range of I/O operations found in ML workloads. Currently, I/O analysis tools often fall short in a number of ways, including the inability to identify I/O requests at the thread level and only capturing traces of specific I/O interfaces (i.e., POSIX and STDIO). This makes it challenging to find a general-purpose tool applicable to a wide variety of workloads, which can reduce the consistency in which I/O traces are captured. In addition, many tools (i.e., Darshan (et al, 2016c) and Scalasca (Geimer et al., 2010)) target MPI applications limiting their compatibility and effectiveness with other communication libraries (i.e., GLOO (Facebook Incubator, 2024), Spark (Apache Software Foundation, 2024a), and Dask (Dask, 2024)). It should be noted that Darshan is capable of capturing I/O traces from applications that do not use MPI through the NONMPI environment variable. However, this results in the output of a single DXT log for each process instead of the more common single file output that captures all I/O traces from an application using MPI. In some cases, this leaves developers with the task of manually merging the results from their non-MPI applications in order to conduct further analysis.
The topic of ML I/O and whether it has a significant role in poor performance has been a debate within the HPC community. Many point out that HPC systems often have enough node-local storage to cache the dataset, resulting in minimal time spent waiting on I/O (Schimmelpfennig et al., 2021; et al, 2021h, g, 2018d). Others mention that not all HPC systems have large amounts of node local storage or the datasets used are simply too large to be cached (Schimmelpfennig et al., 2021; et al, 2021h; Xu et al., 2021; et al, 2018d). In addition, due to the random sampling that commonly occurs during ML training, modern PFSs can often become an I/O bottleneck (Petersen, 2015; et al, 2020k, 2018d). Ultimately, whether I/O is a significant issue is subjective to the workload at hand. It depends on various factors such as dataset size, storage capacity, storage system bandwidth, I/O library implementation and their configurations, and I/O access patterns of applications. Furthermore, many ML application developers are not HPC storage experts; hence, even if their system is capable of providing great performance, transparent solutions such as caching and prefetching are needed to ensure the HPC systems are utilized effectively. Transparent solutions are of particular interest due to the increasing number of heterogeneous processors used within ML applications. These increase the complexity of optimally utilizing available compute resources. For these reasons, we believe that I/O and data management are key areas of interest to the ML community.
There are several common phases in an ML application lifecycle: data generation, data preparation, training, and inference. The data management found within each phase varies widely due to different I/O access patterns, data sources, and file formats. We provide a taxonomy of the data management found throughout the ML lifecycle in Fig. 3. The ML lifecycle begins with data generation, which involves collecting data from various sources (e.g., simulation results or web scraping), which are then stored using a chosen file format (e.g., .tfrecord, .csv, or .png) and storage location (e.g., SSD, disk, cloud, or database). This is commonly followed by an offline data preparation phase, which performs a series of transformations and reductions (i.e., filters, aggregations, or generalizations) over the collected data to increase the training efficiency and accuracy of the ML model. The training process involves reading samples from the data sources, which are then fed to computations. The order in which the samples are read largely depends on the chosen optimizer (e.g., Stochastic Gradient Descent and Adaptive Moment Estimation or Adam, etc.) and training distribution strategy (e.g., model or data parallelism). Due to many optimizers requiring random sampling and storage systems being optimized for contiguous sequential reading, training I/O is a potential bottleneck. Once the model has been trained, it is often deployed and fed “new” data during the inference phase. It is common for the online and offline data preparation phases to occur before and during the inference phase as well. More details on each phase of ML can be found in Section 3, but it is important to note the wide range of factors that impact I/O and data management throughout the ML lifecycle, such as varying file formats, data sources, data preparation requirements, and data modalities, among others. It is impractical to cover all possible combinations of the factors mentioned above, and therefore, in this paper, we aim to discuss the I/O approaches found in the most common ML applications running on HPC systems.
The structure of the paper is as follows. First, in Section 2, we define the scope of the survey and present trends in the cited literature. Then, in Section 3, we describe the data formats and modalities found in ML workloads and provide an overview of the typical phases of ML applications and the popular methods of distributing the model training on HPC systems. Next, in Section 4, we describe the benchmarks and profiling tools scientists and researchers can use to ensure proper system utilization under ML workloads. In Section 5, we describe the I/O access patterns found in the offline data preparation, training phase, and inference phases. The training phase is exemplified by analyzing the currently available representative I/O benchmarks. We then discuss, in Section 6, various I/O optimization techniques used in modern ML frameworks and those proposed in the recent research literature. Lastly, in Section 7, we discuss the gaps in the current research of ML I/O in HPC systems.
2. Scope of the Survey
ML applications use various environments, from local machines for small-scale applications to HPC and cloud systems for large-scale modeling. This survey aims to provide a comprehensive review of the I/O patterns in ML running on HPC systems. This means that I/O patterns and other data management solutions that may be common in other environments (e.g., cloud or local) are not relevant. For example, in Sec 5, we discuss only the I/O access patterns found in common HPC scenarios and exclude I/O access patterns that may be common in cloud or local environments.
@@ 2. Scope of the Survey
To ensure a comprehensive review of the current literature, we have selected a -year time window from to and used ACM Digital Library and IEEE Xplore to filter by papers that mention the following keywords: “I/O Access Patterns”, “I/O pattern”, “HPC”, “Storage performance”, and “I/O performance in AI/ML applications”. We found additional papers through the citations of the work initially discovered using the digital libraries for a total of papers. The various subjects explored in the papers surveyed are shown in Table 1. We selected papers from leading conferences and journals in HPC and AI areas including:
@@
IEEE Conference on High Performance Extreme Computing (HPEC)
@@
IEEE Modeling, Analysis, and Simulation of Computer and Telecommunication Systems (MASCOTS)
@@
The parallel coordinates diagram in Fig. 2 illustrates the popularity of various software tools, frameworks, topics, and file formats in the surveyed papers. TensorFlow was discussed in papers, while PyTorch appeared in papers, making these two ML frameworks the most popular by a significant margin. Training emerged as the most common area of focus, with a total of papers, followed by checkpointing, with a total of papers, and lastly, inference, with papers. Furthermore, HDF5 and TFRecord were notably mentioned in and of the surveyed papers, respectively. Based on these results, there has been a strong focus on optimizing I/O for training ML models using TensorFlow and PyTorch over the last couple of years.
In this paper, our contributions are a presentation and analysis of the I/O patterns found when emulating two renowned ML workloads: BERT (Devlin et al., 2019a), and Unet3D (Özgün Çiçek et al., 2016). In addition, we conduct a review of the currently proposed I/O optimization techniques through R&D efforts and present several areas where future research is needed. Understanding the I/O patterns found in ML applications allows model training speeds to be reduced. This contributes to the rapid development of ML models, which can be used to perform tasks such as medical imaging analysis and help in the creation of autonomous vehicles. It should be noted that throughout the survey, we use the terms “AI” and “ML” synonymously, as is commonly done in the industry. However, “ML” is the more precise term referring to the concept in which machines extract knowledge from data and learn from it.
3. ML Data: Data Formats, Modalities, and Common Phases
In this section, we describe the ML lifecycle and the data formats and modalities commonly found throughout this lifecycle. We also discuss the common phases ML applications undergo and the ways of distributing these phases to exploit the resources offered by HPC systems. We then discuss model checkpointing and its impact on training speeds.
ML applications typically undergo four phases during the ML lifecycle: data generation, dataset preparation, training, and inference (et al, 2021e). We show a taxonomy of the data management operations and characteristics throughout this lifecycle in Fig. 3. During data generation, data is obtained through various sources and methods such as simulations, web scraping, and public datasets. During this phase, properties such as the dataset modality, file format, and data storage location are often decided. The data preparation phase is divided into various data preparation steps, including transformation, augmentation, and splitting. The training arm of the taxonomy consists of data I/O operations across common training algorithms and checkpointing tasks. Inference includes a taxonomy similar to that of the training phase. As can be seen, data is a prominent component throughout the entirety of the ML lifecycle.
@@ 3.1. Data Formats and Dataset Modalities
Data is at the forefront of the ML lifecycle, and without the correct quantity and quality of data, machine learning models fail to achieve optimal performance and generalization, disabling their ability to provide accurate insights and solutions. Here, we describe the file formats and data modalities commonly used during model training and inference. The file formats described were chosen due to their frequent appearance in the cited literature or their connection with state-of-the-art ML training and preprocessing frameworks (e.g., .tfrecord – TensorFlow and .parquet – Apache Parquet).
@@
| LibriSpeech (Panayotov et al., 2015) | GiB | .flac, .txt | ||
@@
| UCR Time Series Archive (Dau, 2018) | MiB | .tsv | ||
@@
| Text | refinedweb (et al, 2023d) | TiB | .parquet | |
@@
| The Stack v2 (et al, 2024) | TiB | Code Files |
The selection of file formats model developers can choose from is often constrained by the data modality in use. The dataset modality refers to the mode or type of data being stored (e.g., audio, video, image, etc.). Due to the large variety of ML model goals (e.g., image recognition, optical character recognition (OCR), relationship detection, etc.), modalities vary widely. The dataset modality has a significant impact on the viable file formats available and, therefore, the common I/O operations required to load and use the dataset. In Table 3, we list the dataset modalities commonly used for ML applications and three corresponding publicly available datasets often used for ML research, exemplifying each modality. The provided values are approximate and subject to variation depending on which version of the dataset is selected. For instance, the Charades (Sigurdsson et al., 2016) dataset is available at different video resolutions, while the CIFAR-10 (Krizhevsky, 2009) dataset is available as a custom binary format or as a binary containing a Python “Pickled” object. Additionally, in certain cases, like AudioSet, links to the video data are provided rather than the raw video data itself, allowing users to choose the preferred video format.